OpenStudy (anonymous):

Which of the following is not a subset of the set of rational numbers? (Points : 1) Natural Numbers Integers Whole Numbers Real Numbers

OpenStudy (amistre64):

what defines a number as ratioanl?

OpenStudy (anonymous):

Rational Number Can Be Written As A Fraction

OpenStudy (anonymous):

Note that \(X \subset Y\) if for every \(x \in X\) we have \(x \in Y\).

OpenStudy (amistre64):

"as a fraction" is close ... but it requires a little more care.\[\frac{\sqrt{2}}{3}\]is written like a fraction, but it is not a ratioanl number

OpenStudy (amistre64):

the top and bottom of the "fraction" have to be expressed as integers

OpenStudy (anonymous):

The precise definition is any number of the form \(\frac{p}{q}\) with \(p, q \in \mathbb{Z}\) and \(q \neq 0\)

OpenStudy (amistre64):

the top 3 options all consist of integers, and can be written in a rational form. can you think of a real number that cannot be written "like a fraction"?

OpenStudy (anonymous):

Is there an a number \(x\) in the set of real numbers which is not rational? Is there an \(x \in \mathbb{R}\) such that \(x \not\in \mathbb{Q}\)?